9/10 times 4/15
Find common denominators then simplify.

Answer by lenny460(1073)

9/10 times 4/15
Find common denominators then simplify.
9/10 * 4/15
9 * 4 = 36
10 * 15 = 150
36/ 150 = 18/75 = 6/25

Answer:
6/25

Decimal Form:
0.24